OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (moongazer):

-1

OpenStudy (anonymous):

y=x-3 x+3(x-3)=5 x+3x-9=5 4x-9=5 4x=14 x=3.5 then solve for y

OpenStudy (moongazer):

3x-3y=-9 x+3y =5 add it for the x 4x=-4 x=-1

OpenStudy (moongazer):

@tim538 That is wrong. y should be equal to "y= x+3" not "y= x-3"

OpenStudy (anonymous):

sorry your right

OpenStudy (moongazer):

:)